Calculate the degrees of freedom for a difference between means test, where one sample is N=72 and the other is N=55
Answer #1

Answer)

There are two scenarios

If it is mentioned that population variances are equal then

Degrees of freedom is given by n1 + n2 - 2 = 72 + 55 - 2 = 125

If it is mentioned that population variances are unequal then

Degrees of freedom is = smaller of n1-1 and n2-1 = 54